### v1

# Measurements of $K_{S}^{0}$-$K_{L}^{0}$ asymmetries in the decays $\Lambda_{c}^{+}\to pK_{L,S}^{0}$, $pK_{L,S}^{0}\pi^{+}\pi^{-}$ and $pK_{L,S}^{0}\pi^{0}$

Abstract
Using $e^{+}e^{-}$ annihilation data sets corresponding to an integrated luminosity of 4.5 $\text{fb}^{-1}$, collected with the BESIII detector at center-of-mass energies between 4.600 and 4.699 GeV, we report the first measurements of the absolute branching fractions $\mathcal{B}(\Lambda_{c}^{+}\to pK_{L}^{0})=(1.67\pm 0.06\pm 0.04)\%$, $\mathcal{B}(\Lambda_{c}^{+}\to pK_{L}^{0}\pi^{+}\pi^{-})=(1.69\pm 0.10\pm 0.05)\%$, and $\mathcal{B}(\Lambda_{c}^{+}\to pK_{L}^{0}\pi^{0})=(2.02\pm 0.13\pm 0.05)\%$, where the first uncertainties are statistical and the second systematic. Combining with the known branching fractions of $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda_{c}^{+}\to pK_{S}^{0}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK_{S}^{0}\pi^{0}$, we present the first measurements of the $K_{S}^{0}$-$K_{L}^{0}$ asymmetries $R(\Lambda_{c}^{+},K_{S,L}^{0}X)=\frac{\mathcal{B}(\Lambda_{c}^{+}\to K_{S}^{0}X)-\mathcal{B}(\Lambda_{c}^{+}\to K_{L}^{0}X)}{\mathcal{B}(\Lambda_{c}^{+}\to K_{S}^{0}X)+\mathcal{B}(\Lambda_{c}^{+}\to K_{L}^{0}X)}$ in charmed baryon decays: $R(\Lambda_{c}^{+},pK_{S,L}^{0})=-0.025\pm 0.031$, $R(\Lambda_{c}^{+},pK_{S,L}^{0}\pi^{+}\pi^{-})=-0.027\pm 0.048$, and $R(\Lambda_{c}^{+},pK_{S,L}^{0}\pi^{0})=-0.015\pm 0.046$. No significant asymmetries within the uncertainties are observed.
Keywords:
$\Lambda_{c}^{+}$ baryon, $K^{0}_{S}$-$K^{0}_{L}$ asymmetry, the BESIII detector
arxiv: 1234.56789
1 Introduction
The lightest charmed baryon, $\Lambda_{c}^{+}$, provides a unique environment for studying the behavior of light di-quarks in the presence of a heavy quark Cheng:2021qpd . Its hadronic decays occur only through the weak interaction, and various theoretical models have been proposed. These include the covariant confined quark model Korner:1992wi ; Ivanov:1997ra , the pole model Cheng:1991sn ; Xu:1992vc ; Cheng:1993gf ; Xu:1992sw ; Zenczykowski:1993jm ; Sharma:1998rd , current algebra Korner:1978ec ; Uppal:1994pt , and SU(3) flavor symmetry approaches Lu:2016ogy ; Savage:1989qr ; Kohara:1991ug ; Verma:1995dk ; Chau:1995gk . Its decay falls into three categories: Cabibbo-favored (CF) decays, singly Cabibbo-suppressed decays, and doubly Cabibbo-suppressed (DCS) decays. The decay amplitudes of the CF and DCS modes are expected to be proportional to the products of the Cabibbo-Kobayashi-Maskawa elements $|V_{ud}^{*}V_{cs}|$ and $|V_{us}^{*}V_{cd}|$, respectively. The ratio of their decays is approximately of the order of $\mathcal{O}(10^{-3})$, resulting in a small branching fraction (BF) for the DCS decay and making it challenging to observe directly in experiments.
In addition to direct measurements of DCS decays, the amplitudes of DCS modes can be probed using the $K^{0}_{S}$-$K^{0}_{L}$ asymmetry in the decays into neutral kaons, which arises from the interference between CF and DCS amplitudes Bigi:1994aw ; Wang:2017gxe . The $K^{0}_{S}$-$K^{0}_{L}$ asymmetry has been studied in the decays of charmed $D$ mesons, where the asymmetry is defined by
$R(D,K_{S,L}^{0}X)=\frac{\mathcal{B}(D\to K^{0}_{S}X)-\mathcal{B}(D\to K^{0}_{L}X)}{\mathcal{B}(D\to K^{0}_{S}X)+\mathcal{B}(D\to K^{0}_{L}X)},$ (1)
and $X$ can be $\pi^{0}$, $\eta$, $\eta^{\prime}$, $\omega$, $\rho^{0}$ or $\phi$. A large asymmetry for $R(D^{0},K_{S,L}^{0}\pi^{0})$ was reported in a previous measurement by the CLEO experiment as $R(D^{0},K_{S,L}^{0}\pi^{0})=0.108\pm 0.025\pm 0.024$ CLEO:2007rhw , where the first uncertainty is statistical and the second is systematic. The BESIII experiment reported measurements of the $K^{0}_{S}$-$K^{0}_{L}$ asymmetries $R(D^{0},K_{S,L}^{0}X)$, where $X=\phi$, $\eta$, $\omega$ BESIII:2022xhe . Significant asymmetries were observed in $D^{0}\to K^{0}_{L}\eta$ and $D^{0}\to K^{0}_{L}\eta^{\prime}$ decays with $R(D^{0},K_{S,L}^{0}\eta)=0.080\pm 0.022$ and $R(D^{0},K_{S,L}^{0}\eta^{\prime})=0.108\pm 0.035$, respectively. In addition, this asymmetry has been investigated for the lightest charmed strange meson, and $R(D_{s}^{+},K_{S,L}^{0}K^{+})$ was determined to be $(-2.1\pm 1.9\pm 1.6)\%$ BESIII:2019kfh . However, such measurements have not been made for the decays of charmed baryons.
Using flavor SU(3) asymmetry Lu:2016ogy ; Savage:1989qr ; Kohara:1991ug ; Verma:1995dk ; Chau:1995gk , theoretical predictions Wang:2017gxe for $K^{0}_{S}$-$K^{0}_{L}$ asymmetries have been made for charmed baryon two-body decays into a light baryon and a neutral kaon. Similar to Equation 1, the asymmetry of $\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{S}X)$ and $\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{L}X)$ in charmed baryon decays is defined as
$R(\Lambda_{c}^{+},K_{S,L}^{0}X)=\frac{\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{S}X)-\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{L}X)}{\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{S}X)+\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{L}X)},$ (2)
where $X$ is $p$, $p\pi^{+}\pi^{-}$ or $p\pi^{0}$. Equation 2 can be further reduced as $R(\Lambda_{c}^{+}\to K^{0}_{S,L}X)\simeq-2r_{f}\cos{\delta_{f}}$, where $r_{f}$ and $\delta_{f}$ are the relative strength and phase between the DCS ($\Lambda_{c}^{+}\to K^{0}X$) and CF ($\Lambda_{c}^{+}\to\bar{K}^{0}X$) amplitudes, respectively. The parameter $r_{f}$ is expected to be proportional to the ratio $|V_{cd}^{*}V_{us}/V_{cs}^{*}V_{ud}|\sim\lambda^{2}$ ParticleDataGroup:2022pth . A non-zero asymmetry value indicates the presence of DCS processes. The asymmetry of $\Lambda_{c}^{+}\to pK_{S,L}^{0}$ is predicted to be in the range of ($-0.010$, $0.087$) in Ref. Wang:2017gxe . The $K^{0}_{S}$-$K^{0}_{L}$ asymmetry is a promising observable with which to search for the two-body DCS processes of charmed baryons.
In this paper, we report the first measurements of the absolute BFs of $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ based on $e^{+}e^{-}$ annihilation data samples corresponding to a total integrated luminosity of 4.5 fb-1 collected at the center-of-mass (c.m.) energies $\sqrt{s}$ between 4.600 and 4.699 GeV. The luminosities are listed in Table 1 BESIII:2022dxl ; BESIII:2022ulv . Using the results of $\mathcal{B}(\Lambda_{c}^{+}\to pK^{0}_{S})$, $\mathcal{B}(\Lambda_{c}^{+}\to pK^{0}_{S}\pi^{+}\pi^{-})$, and $\mathcal{B}(\Lambda_{c}^{+}\to pK^{0}_{S}\pi^{0})$ from the Particle Data Group (PDG) ParticleDataGroup:2022pth , we present the $K^{0}_{S}$-$K^{0}_{L}$ asymmetries $R(\Lambda_{c}^{+},K_{S,L}^{0}X)$, where $X=p$, $p\pi^{+}\pi^{-}$ or $p\pi^{0}$. Charge conjugate channels are implied throughout this paper, unless explicitly stated.
Table 1: The integrated luminosities at each c.m. energy BESIII:2022dxl ; BESIII:2022ulv .
[…]
4.600 586.9 $\pm$ 0.1 $\pm$ 3.9
4.612 103.7 $\pm$ 0.1 $\pm$ 0.6
4.628 521.5 $\pm$ 0.1 $\pm$ 2.8
4.641 551.7 $\pm$ 0.1 $\pm$ 2.9
4.661 529.4 $\pm$ 0.1 $\pm$ 2.8
4.682 1667.4 $\pm$ 0.2 $\pm$ 8.8
4.699 535.5 $\pm$ 0.1 $\pm$ 2.8
2 BESIII experiment and Monte Carlo simulation
[…]
Simulated samples generated with geant4-based GEANT4:2002zbu Monte Carlo (MC) software, which includes the geometric description of the BESIII detector and the detector response performance YouZhengYun2008 ; Liang:2009zzb ; Huang:2022wuo , are used to determine detection efficiencies and to estimate potential background contributions. The simulation describes the beam energy spread and the initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc Jadach:2000ir ; Jadach:1999vf . The inclusive MC samples, corresponding to about 40 times the number of events of the data samples, include the production of $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ pairs, open charm processes, the ISR production of vector charmonium(-like) states, and the continuum processes incorporated in kkmc Jadach:2000ir ; Jadach:1999vf . The known decay modes are modeled with evtgen Lange:2001uf ; Ping:2008zz using BFs taken from the PDG ParticleDataGroup:2022pth , and the remaining unknown charmonium decays are modeled with lundcharm Chen:2000tv ; Lange:2001uf . Final state radiation from charged final state particles is incorporated using photos Richter-Was:1992hxq . For the production of $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ events, the Born cross-section line shape from BESIII measurements is used BESIII:2017kqg ; BESIII:2023rwv . Exclusive $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ signal MC samples are generated with $\bar{\Lambda}_{c}^{-}$ decaying to twelve specific tag modes (as described in Section 3) and $\Lambda_{c}^{+}$ decaying to $pK^{0}_{L}$, $pK^{0}_{L}\pi^{+}\pi^{-}$ and $pK^{0}_{L}\pi^{0}$. The angular distribution of the decay $\Lambda_{c}^{+}\to pK^{0}_{L}$ is modeled with decay asymmetry parameters obtained from Ref. BESIII:2019odb . For processes from $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ channels, signal models are tunned based on the data. Additional MC samples are generated to estimate contributions from peaking background processes, where $\bar{\Lambda}_{c}^{-}$ decays into tag modes and $\Lambda_{c}^{+}$ decays into $pK^{0}_{S}$, $p\eta$, $pK^{0}_{S}\pi^{0}$, and $pK^{0}_{S}\pi^{+}\pi^{-}$, with $K^{0}_{S}$ and $\eta$ decaying inclusively. Each tag mode of the exclusive MC samples is generated with the same number of events.
3 Data analysis
Taking advantage of the threshold production of the $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ pair, the double-tag (DT) method MARK-III:1985hbd ; MARK-III:1987jsm ; Ke:2023qzc ; Li:2021iwf is employed to study $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, where $K^{0}_{L}$ is reconstructed by the missing-mass technique. A single-tag (ST) event is selected by tagging a $\bar{\Lambda}_{c}^{-}$ baryon with one of the following twelve tag modes: $\bar{p}K_{S}^{0}$, $\bar{p}K^{+}\pi^{-}$, $\bar{p}K_{S}^{0}\pi^{0}$, $\bar{p}K_{S}^{0}\pi^{-}\pi^{+}$, $\bar{p}K^{+}\pi^{-}\pi^{0}$, $\bar{p}\pi^{-}\pi^{+}$, $\bar{\Lambda}\pi^{-}$, $\bar{\Lambda}\pi^{-}\pi^{0}$, $\bar{\Lambda}\pi^{-}\pi^{+}\pi^{-}$, $\bar{\Sigma}^{0}\pi^{-}$, $\bar{\Sigma}^{-}\pi^{0}$, and $\bar{\Sigma}^{-}\pi^{-}\pi^{+}$. The ST event selection criteria, efficiencies, and yields are described in Ref. BESIII:2022xne . The signal decays $\Lambda_{c}^{+}\to pK^{0}_{L}$, $pK^{0}_{L}\pi^{+}\pi^{-}$, and $pK^{0}_{L}\pi^{0}$ are reconstructed using the remaining charged tracks and photons recoiling against the ST $\bar{\Lambda}_{c}^{-}$ candidates, and referred to as DT events.
[…]
Photon candidates are reconstructed from showers that are not associated with any charged tracks in the EMC BESIII:2009fln . The deposited energy of each shower in the EMC is required to be greater than 25 MeV in the barrel region ($|\cos\theta|<0.80$), and greater than 50 MeV in the end cap region ($0.920.86<|\cos\theta|<$ 0.92). The EMC time difference from the event start time is required to be less than 700 ns, to exclude electronic noise and showers unrelated to the events. The opening angle between each shower and $\bar{p}$ must be greater than $20^{\circ}$, to suppress the background from annihilation of $\bar{p}$ with the detector material. The $\pi^{0}$ candidates are reconstructed from photon pairs with invariant mass $M(\gamma\gamma)$ in the range 0.115 GeV/$c^{2}$ $<M(\gamma\gamma)<$ 0.150 GeV/$c^{2}$. To improve momentum resolution and exclude background, a kinematic fit is performed to constrain $M(\gamma\gamma)$ to the known $\pi^{0}$ mass ParticleDataGroup:2022pth , and candidates with fit quality $\chi^{2}<20$ are retained for further analysis.
The signal candidates of $\Lambda_{c}^{+}\to pK^{0}_{L}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ are required to have only one charged track with opposite charge to the tagged $\bar{\Lambda}_{c}^{-}$ satisfying the proton PID criteria. For $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ decay, the $\pi^{0}$ candidate with the highest energy is selected. In the reconstruction of $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, events must have only three remaining charged tracks with correct charges and PID. Candidates with additional charged tracks, whose distances of closest approaches to the IP are within $\pm$20 cm along the beam direction, are excluded. The presence of the $K^{0}_{L}$ is inferred by the kinematic variable $M^{2}_{\mathrm{miss}}$, defined as
$M^{2}_{\mathrm{miss}}\equiv\left(E_{\mathrm{beam}}-E_{\mathrm{selected}}\right)^{2}/c^{4}-\left|\vec{p}_{\Lambda_{c}^{+}}-\vec{p}_{\mathrm{selected}}\right|^{2}/c^{2},$ (3)
where $E_{\mathrm{beam}}$ is the beam energy and $E_{\mathrm{selected}}$ ($\vec{p}_{\mathrm{selected}}$) is the total measured energy (momentum) of the selected particles in the DT signal side, boosted into the c.m. system of $e^{+}e^{-}$. To improve the momentum resolution, the momentum of $\Lambda_{c}^{+}$ is determined by
$\vec{p}_{\Lambda_{c}^{+}}\equiv-\hat{p}_{\mathrm{\bar{\Lambda}_{c}^{-}}}\sqrt{E_{\mathrm{beam}}^{2}/c^{2}-m^{2}_{\Lambda_{c}^{+}}c^{2}},$ (4)
where $\hat{p}_{\mathrm{\bar{\Lambda}_{c}^{-}}}$ is the direction of the tagged $\bar{\Lambda}_{c}^{-}$ and $m_{\Lambda_{c}^{+}}$ is the known $\Lambda_{c}^{+}$ baryon mass taken from the PDG ParticleDataGroup:2022pth . For all three decays, the $M^{2}_{\mathrm{miss}}$ distributions are expected to have a peak around the known mass squared of $K^{0}_{L}$ ParticleDataGroup:2022pth .
[…]
To improve the momentum resolution, a six constraint (6C) kinematic fit is performed requiring total four-momentum conservation with respect to that of the initial $e^{+}e^{-}$ collision and constraining both masses of the tagged $\bar{\Lambda}_{c}^{-}$ and the signal $\Lambda_{c}^{+}$ to $m_{\Lambda_{c}^{+}}$. The $K^{0}_{L}$ is treated as a missing particle, and its four-momentum and mass are free in the kinematic fit. The $\chi^{2}$ of the kinematic fit for each signal mode is required to be less than the optimized value that maximizes the figure of merit $S/\sqrt{S+B}$, where $S$ and $B$ are the numbers of signal and background events from MC simulations, scaled to the data luminosity. The optimized requirements are $\chi^{2}<60$ for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\chi^{2}<25$ for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\chi^{2}<20$ for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$. The resulting $M^{2}_{\mathrm{miss}}$ distributions of the DT events are shown in Figure 1, which combine all data samples at the seven c.m. energies. Signal events are indicated by the significant peaks around the $K^{0}_{L}$ mass squared.
Figure 1: The $M^{2}_{\mathrm{miss}}$ distributions of the selected DT events for (a) $\Lambda_{c}^{+}\to pK^{0}_{L}$, (b) $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and (c) $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ decays. The points with error bars are data combined from seven c.m. energies, the red histograms indicate the signal processes, and the green and orange histograms are the peaking backgrounds. The cyan and violet histograms represent non-signal $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ and inclusive hadronic background processes, respectively.
There are peaking backgrounds remaining from $\Lambda_{c}^{+}\to pK^{0}_{S}(\to\pi^{0}\pi^{0})$ and $\Lambda_{c}^{+}\to p\eta(\to\gamma\gamma$ or $3\pi^{0})$, $\Lambda_{c}^{+}\to pK^{0}_{S}(\to\pi^{0}\pi^{0})\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK^{0}_{S}(\to\pi^{0}\pi^{0})\pi^{0}$ in the corresponding signal modes. The peaking background events from $\Lambda_{c}^{+}\to K^{0}_{S}X$ decays $N^{\mathrm{Bkg}}_{K^{0}_{S}X}$ are determined by
$N^{\mathrm{Bkg}}_{K^{0}_{S}X}=N^{\mathrm{Data}}_{\mathrm{DT}\,K^{0}_{S}X}\cdot w_{K^{0}_{S}X},\ w_{K^{0}_{S}X}=\frac{\sum_{i}s_{i}\cdot\frac{N^{\mathrm{ST}}_{i}}{\varepsilon^{\mathrm{ST}}_{i}}\cdot N^{\mathrm{MC},i}_{\mathrm{DT}\,K^{0}_{L}X}}{\sum_{i}\frac{N^{\mathrm{ST}}_{i}}{\varepsilon^{\mathrm{ST}}_{i}}\cdot N^{\mathrm{MC},i}_{\mathrm{DT}\,K^{0}_{S}X}},$ (5)
where $i$ represents the tag mode, and $N^{\mathrm{Data}}_{\mathrm{DT}\,K^{0}_{S}X}$ denotes the data yields passing the DT selection criteria of $\Lambda_{c}^{+}\to K^{0}_{S}X$. Here, the DT selection criteria of $\Lambda_{c}^{+}\to K^{0}_{S}X$ require a fully reconstructed $K^{0}_{S}$ from $\pi^{+}\pi^{-}$ combinations, as described in Ref. BESIII:2015bjk . $N^{\mathrm{Data}}_{\mathrm{DT}\,K^{0}_{S}X}$ is corrected by the factor $w_{K^{0}_{S}X}$, which is derived from the exclusive MC simulation samples of $\Lambda_{c}^{+}\to K^{0}_{S}X$. $N_{\mathrm{DT}\,K^{0}_{L}X}^{\mathrm{MC},i}$ and $N_{\mathrm{DT}\,K^{0}_{S}X}^{\mathrm{MC},i}$ are the numbers of the $K^{0}_{S}X$ MC events that satisfy the DT selection criteria of $\Lambda_{c}^{+}\to K^{0}_{L}X$ and $\Lambda_{c}^{+}\to K^{0}_{S}X$, respectively. $N^{\mathrm{ST}}_{i}$ and $\varepsilon^{\mathrm{ST}}_{i}$ are the ST yields and ST efficiencies from Ref. BESIII:2022xne . A scale factor $s_{i}$ is specified for each tag mode, and $s_{i}$ is set to 2 if both the tag and signal modes are $K^{0}_{S}X$. Otherwise, it is set to 1. For peaking background events from $\Lambda_{c}^{+}\to p\eta$, the contribution is evaluated based on the corresponding exclusive MC samples using
$N^{\mathrm{Bkg}}_{p\eta}=\mathcal{B}(\Lambda_{c}^{+}\to p\eta)\cdot w_{p\eta},\ w_{p\eta}=\sum_{i}\left(\frac{N^{\mathrm{ST}}_{i}}{\varepsilon^{\mathrm{ST}}_{i}}\cdot\frac{N^{\prime\mathrm{MC},i}_{pK^{0}_{L}}}{N^{\prime\mathrm{MC},i}_{\mathrm{tot}}}\right),$ (6)
with $\mathcal{B}(\Lambda_{c}^{+}\to p\eta)=(1.41\pm 0.11)\times 10^{-3}$ ParticleDataGroup:2022pth . $N^{\prime\mathrm{MC},i}_{pK^{0}_{L}}$ is the number of surviving DT events for the $i$-th tag mode, that satisfy the DT selection criteria of $\Lambda_{c}^{+}\to pK^{0}_{L}$, and $N^{\prime\mathrm{MC},i}_{\mathrm{tot}}$ is the total number of MC events generated for the $i$-th tag mode. Table 2 summarizes the contributions arising from each peaking background process.
Table 2: Estimated yields of peaking backgrounds at each c.m. energy. The uncertainties are statistical only.
$\sqrt{s}$ (GeV) $\Lambda_{c}^{+}\to pK^{0}_{S}$ $\Lambda_{c}^{+}\to p\eta$ $\Lambda_{c}^{+}\to pK^{0}_{S}\pi^{+}\pi^{-}$ $\Lambda_{c}^{+}\to pK^{0}_{S}\pi^{0}$
4.600 59 $\pm$ 6 13 $\pm$ 1 25 $\pm$ 5 44 $\pm$ 6
4.612 14 $\pm$ 3 2.3$\pm$ 0.2 4 $\pm$ 1 8 $\pm$ 3
4.628 70 $\pm$ 7 11 $\pm$ 1 31 $\pm$ 4 38 $\pm$ 6
4.641 68 $\pm$ 7 12 $\pm$ 1 30 $\pm$ 5 38 $\pm$ 7
4.661 60 $\pm$ 6 12 $\pm$ 1 39 $\pm$ 5 51 $\pm$ 8
4.682 198 $\pm$ 12 35 $\pm$ 3 97 $\pm$ 9 150 $\pm$ 13
4.699 54 $\pm$ 6 10 $\pm$ 1 24 $\pm$ 5 37 $\pm$ 6
A simultaneous unbinned maximum-likelihood fit is performed on the $M^{2}_{\mathrm{miss}}$ distributions of the seven c.m. energies. The signal and peaking backgrounds are modeled by individual MC-simulated shapes convolved with Gaussian functions to account for differences between the data and MC simulations. The Gaussian means and widths are free parameters in the fit. The yields of the peaking background events are free with their mean and standard deviation values set to the results listed in Table 2. For the signal mode $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, a truth-matching method is employed to obtain the pure signal shape by comparing the two photons from the $\pi^{0}$ with their corresponding MC truth information. The opening angle $\theta_{\mathrm{truth}}$ between the truth and the reconstructed photons is required to be less than $10^{\circ}$. The combinatorial background shape is taken from the inclusive MC samples, including non-signal $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ and continuum hadron production events.
The BFs of the decays $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ are shared variables for the seven c.m. energies in the simultaneous fit, determined by
$\mathcal{B}_{\mathrm{sig}}=\frac{N^{\mathrm{DT}}}{N^{\mathrm{ST}}\cdot\varepsilon_{\mathrm{avg}}\cdot\mathcal{B}_{\mathrm{int}}},$ (7)
where $\varepsilon_{\mathrm{avg}}=\left(\sum_{i}N^{\mathrm{ST}}_{i}\cdot\varepsilon^{\mathrm{DT}}_{i}/\varepsilon^{\mathrm{ST}}_{i}\right)/N^{\mathrm{ST}}$ is the average detection efficiency for detecting signal modes in ST events and $i$ represents the $i$-th ST tag mode. Table 3 lists the ST events and the average detection efficiencies for each c.m. energy. $N^{\mathrm{DT}}$ and $\varepsilon^{\mathrm{DT}}_{i}$ are the DT yields and corresponding efficiencies, respectively. $\mathcal{B}_{\mathrm{int}}$ is the intermediate BF of $\pi^{0}$, $\mathcal{B}(\pi^{0}\to\gamma\gamma)=(98.823\pm 0.034)\%$ ParticleDataGroup:2022pth for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ decay. Figure 2 shows the results of fits to the $M^{2}_{\mathrm{miss}}$ distributions, combining all data samples. From these fits, the BFs are $\mathcal{B}(\Lambda_{c}^{+}\to pK^{0}_{L})=(1.67\pm 0.06)\%$, $\mathcal{B}(\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-})=(1.69\pm 0.10)\%$, and $\mathcal{B}(\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0})=(2.02\pm 0.13)\%$, where the uncertainties are statistical only. The total DT signal yields from all c.m. energies are $N^{\mathrm{DT}}_{pK^{0}_{L}}=1627\pm 56$, $N^{\mathrm{DT}}_{pK^{0}_{L}\pi^{+}\pi^{-}}=648\pm 39$, and $N^{\mathrm{DT}}_{pK^{0}_{L}\pi^{0}}=652\pm 41$, for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
Table 3: ST events ($N^{\mathrm{ST}}$), average detection efficiencies ($\varepsilon_{\mathrm{avg}}$) and DT signal yields ($N^{\mathrm{DT}}$) for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ decays at each c.m. energy. The errors are statistical only.
[…]
$N^{\mathrm{ST}}$ 17391 $\pm$ 171 3114 $\pm$ 75 14558 $\pm$ 135 15545 $\pm$ 165 15235 $\pm$ 164 44704 $\pm$ 284 12971 $\pm$ 158
modes $\varepsilon_{\mathrm{avg}}(\%)$
$\Lambda_{c}^{+}\to pK^{0}_{L}$ 76.32 $\pm$ 0.08 76.13 $\pm$ 0.08 78.12 $\pm$ 0.08 78.83 $\pm$ 0.08 78.78 $\pm$ 0.08 79.66 $\pm$ 0.09 79.89 $\pm$ 0.09
$\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ 28.61 $\pm$ 0.16 28.86 $\pm$ 0.17 29.87 $\pm$ 0.18 30.96 $\pm$ 0.17 31.12 $\pm$ 0.17 31.93 $\pm$ 0.17 32.83 $\pm$ 0.18
$\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ 23.83 $\pm$ 0.15 24.28 $\pm$ 0.15 25.79 $\pm$ 0.16 26.40 $\pm$ 0.16 26.68 $\pm$ 0.16 27.43 $\pm$ 0.17 27.27 $\pm$ 0.17
modes $N^{\mathrm{DT}}$
$\Lambda_{c}^{+}\to pK^{0}_{L}$ 222 $\pm$ 8 40 $\pm$ 1 190 $\pm$ 7 205 $\pm$ 7 201 $\pm$ 7 596 $\pm$ 21 173 $\pm$ 6
$\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ 84 $\pm$ 5 15 $\pm$ 1 74 $\pm$ 4 81 $\pm$ 5 80 $\pm$ 5 242 $\pm$ 14 72 $\pm$ 4
$\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ 83 $\pm$ 5 15 $\pm$ 1 75 $\pm$ 5 82 $\pm$ 5 81 $\pm$ 5 245 $\pm$ 16 71 $\pm$ 4
Figure 2: Combined fit results of the $M^{2}_{\mathrm{miss}}$ distributions from all data samples for (a) $\Lambda_{c}^{+}\to pK^{0}_{L}$, (b) $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and (c) $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ decays. The black dots with error bars are data, while the black solid curves are the fit results. The red dashed curves are the signal shapes, and the blue and green dashed curves represent the peaking backgrounds $\Lambda_{c}^{+}\to pK^{0}_{S}$ and $\Lambda_{c}^{+}\to p\eta$, respectively. The violet dashed curves are the combinatorial background shapes.
4 Systematic uncertainties
In the DT method, most of the systematic uncertainties associated with the ST selections cancel. The major sources of systematic uncertainties in the BFs measurements are described below and are reported relative to the measured BFs.
  • Tracking and PID efficiencies. The tracking and PID efficiencies of the charged protons and pions are studied using a control sample of $J/\psi\to p\bar{p}\pi^{+}\pi^{-}$ BESIII:2022bkj . The MC simulation samples are weighted by the efficiency ratio between data and MC as function of charged particle momentum and $\cos{\theta}$. The systematic uncertainties of tracking and PID are 0.5% and 0.1% for $\Lambda_{c}^{+}\to pK^{0}_{L}$, 1.6% and 0.8% for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and 0.7% and 0.4% for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
  • No extra charged track requirement. The number of good charged tracks is required to be exactly one (three) for $pK^{0}_{L}$ and $pK^{0}_{L}\pi^{0}$ ($pK^{0}_{L}\pi^{+}\pi^{-}$) DT candidates in the recoil system of the tagged $\bar{\Lambda}_{c}^{-}$. The difference between data and MC simulation from this selection is studied using a control sample of $\Lambda_{c}^{+}\to pK^{+}\pi^{-}$. The systematic uncertainty is 1.9%.
  • MC statistics. The exclusive MC simulation samples are used to obtain the ST and DT detection efficiencies and to estimate the peaking background events. The systematic uncertainties associated with the limited MC sample sizes are estimated to be 0.1%, 0.5%, and 0.5% for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
  • ST yield. The systematic uncertainty arising from the total ST yield is assigned to be 0.2% BESIII:2022xne .
  • Kinematic fit. The model of the MC simulation is much simpler than the real detector performance, resulting in a difference between the data and MC simulation in the track parameters of the charged tracks BESIII:2012mpj . The helix parameters of the charged tracks are corrected, and the BFs are re-evaluated with the updated MC simulation samples. The differences from the measured BFs are taken as the systematic uncertainties associated with the kinematic fit, which are 0.5%, 1.0%, and 0.5% for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
  • Angle($\gamma$,$\bar{p}$) requirement. To estimate the systematic uncertainty of the Angle($\gamma,\bar{p}$) requirement, the difference between the data and MC simulation samples of this requirement is investigated from a control sample of $\psi(3686)\to\pi^{+}\pi^{-}J/\psi,J/\psi\to p\bar{p}\pi^{0}$. The systematic uncertainty is 0.2% for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$.
  • $\pi^{0}$ reconstruction. The systematic uncertainty due to the $\pi^{0}$ reconstruction is determined using the control sample of $J/\psi\to p\bar{p}\pi^{0}$ BESIII:2022ahw . The MC simulation samples are corrected depending on the $\pi^{0}$ momentum. The systematic uncertainty is determined to be 0.5%.
  • Truth-match method. The systematic uncertainty from the truth-match method is determined comparing the measured BFs with and without the truth-match requirements. The resulting systematic uncertainty is taken as 0.2%.
  • Signal model. For $\Lambda_{c}^{+}\to pK^{0}_{L}$, the systematic uncertainty from the signal model is determined varying the decay asymmetry parameters within ±$\pm 1\sigma$. The deviation from the measured BF is found to be negligible. For $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, the signal models in the nominal analysis is tunned based on the data. The possible intermediate resonances are considered in the amplitude analysis, composed of $\Sigma^{*}$, $\Delta^{*}$, $N^{*}$, $\bar{K}^{*}$ and $\rho$. The nominal amplitude models are then replaced by alternative ones with equivalent descriptions of the data. The alternative model of $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ is selected by removing the insignificant intermediate resonances. For $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, the amplitude fit is not stable due to the limited statistics of data. An alternative model is chosen with a similar fit quality to the nominal. The systematic uncertainties are determined to be 1.1% and 0.1% for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
  • Background shape. To investigate the systematic uncertainty from the background shape, the nominal background shape is replaced with a second-order Chebychev polynomial function in the simultaneous fit. The systematic uncertainties are 0.9%, 0.6%, and 0.4% for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
  • Fit bias. The systematic uncertainty from the simultaneous fit is studied with 5000 sets of toy MC samples, which are simulated with all parameters from the fit model fixed. The BFs obtained from the toy samples are fitted with a Gaussian function. The deviations between the Gaussian mean value and nominal BFs are assigned as systematic uncertainties. For the decay $\Lambda_{c}^{+}\to pK_{L}\pi^{0}$, the fit bias is found to be 0.27%, while for the other two signal modes it is negligible.
Other sources of systematic uncertainties, such as the BF of $\pi^{0}\to\gamma\gamma$, are neglected due to their negligible effects. Assuming that all sources of systematic uncertainties in the BFs measurements are uncorrelated, the quadratic sums of the different sources are considered as the total systematic uncertainties, which are 2.2%, 3.1%, and 2.3% for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively. Table 4 lists all the systematic uncertainties discussed above.
Table 4: Relative systematic uncertainties in the BF measurements.
Source $\Lambda_{c}^{+}\to pK^{0}_{L}\ (\%)$ $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}\ (\%)$ $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}\ (\%)$
Tracking 0.5 1.6 0.7
PID 0.1 0.8 0.4
No extra charged track 1.9 1.9 1.9
MC statistics 0.2 0.5 0.5
ST yields 0.2 0.2 0.2
Kinematic fit 0.5 1.0 0.5
Angle($\gamma$,$\bar{p}$) requirement - - 0.2
$\pi^{0}$ reconstruction - - 0.5
Truth-match method - - 0.2
Signal model - 1.1 0.1
Background shape 0.9 0.6 0.4
Fit bias - - 0.3
Total 2.2 3.1 2.3
5 Summary
In summary, we report the BFs of $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ for the first time, by analyzing $e^{+}e^{-}$ annihilation data samples corresponding to an integrated luminosity of 4.5 fb-1 collected at c.m. energies between 4.600 and 4.699 GeV. The measured BFs of these decays are $\mathcal{B}(\Lambda_{c}^{+}\to pK^{0}_{L})=(1.67\pm 0.06\pm 0.04)\%$, $\mathcal{B}(\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-})=(1.69\pm 0.10\pm 0.05)\%$, and $\mathcal{B}(\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0})=(2.02\pm 0.13\pm 0.05)\%$. Combining the BFs measurements in this work with the values of $\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{S}X)$ ParticleDataGroup:2022pth , the $K^{0}_{S}$-$K^{0}_{L}$ asymmetries are determined, as summarized in Table 5. The uncertainties are derived through the standard error propagation procedure, assuming that the uncertainties of the estimated $\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{L}X)$ and the quoted $\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{S}X)$ are uncorrelated. Taking into account the uncertainties, no obvious asymmetry is observed in any of the three decays. The $K^{0}_{S}$-$K^{0}_{L}$ asymmetry of $\Lambda_{c}^{+}\to pK_{S,L}^{0}$ $R(\Lambda_{c}^{+},pK_{S,L}^{0})=-0.025\pm 0.031$ is compatible with the prediction of (--0.010, 0.087) based on SU(3) flavor symmetry Wang:2017gxe . Our measurements of the $K^{0}_{S}$-$K^{0}_{L}$ asymmetries in charmed baryon decays offer the possibility to access the DCS processes involving neutral kaons and provide further constraints on their amplitudes.
Table 5: The BFs $\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{L}X)$, the known BFs $\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{S}X)$, and $K^{0}_{S}$-$K^{0}_{L}$ asymmetries.
Mode $\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{L}X)$ (%) $\mathcal{B}(\Lambda_{c}^{+}\to K^{0}_{S}X)$ (%) ParticleDataGroup:2022pth $R(\Lambda_{c}^{+}$, $K_{L,S}^{0}X)$
$\Lambda_{c}^{+}\to pK_{L,S}^{0}$ 1.67 $\pm$ 0.06 $\pm$ 0.04 1.59 $\pm$ 0.07 -0.025 $\pm$ 0.031
$\Lambda_{c}^{+}\to pK_{L,S}^{0}\pi^{+}\pi^{-}$ 1.69 $\pm$ 0.10 $\pm$ 0.05 1.60 $\pm$ 0.11 -0.027 $\pm$ 0.048
$\Lambda_{c}^{+}\to pK_{L,S}^{0}\pi^{0}$ 2.02 $\pm$ 0.13 $\pm$ 0.05 1.96 $\pm$ 0.12 -0.015 $\pm$ 0.046
Acknowledgements.
      The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2020YFA0406400, 2020YFA0406300, 2023YFA1606000; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11735014, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; German Research Foundation DFG under Contracts Nos. 455635585, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374

### v2

@@
 Combining with the known branching fractions of $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda_{c}^{+}\to pK_{S}^{0}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK_{S}^{0}\pi^{0}$, we present the first measurements of the $K_{S}^{0}$-$K_{L}^{0}$ asymmetries $R(\Lambda_{c}^{+},K_{S,L}^{0}X)=\frac{\mathcal{B}(\Lambda_{c}^{+}\to K_{S}^{0}X)-\mathcal{B}(\Lambda_{c}^{+}\to K_{L}^{0}X)}{\mathcal{B}(\Lambda_{c}^{+}\to K_{S}^{0}X)+\mathcal{B}(\Lambda_{c}^{+}\to K_{L}^{0}X)}$ in charmed baryon decays: $R(\Lambda_{c}^{+},pK_{S,L}^{0})=-0.025\pm 0.031$, $R(\Lambda_{c}^{+},pK_{S,L}^{0}\pi^{+}\pi^{-})=-0.027\pm 0.048$, and $R(\Lambda_{c}^{+},pK_{S,L}^{0}\pi^{0})=-0.015\pm 0.046$.
-No significant asymmetries within the uncertainties are observed.
+No significant asymmetries with statistical significance are observed.
 Keywords:
 $\Lambda_{c}^{+}$ baryon, $K^{0}_{S}$-$K^{0}_{L}$ asymmetry, the BESIII detector
-arxiv: 1234.56789
+arxiv: 2406.18083
 1 Introduction
@@
 These include the covariant confined quark model Korner:1992wi ; Ivanov:1997ra , the pole model Cheng:1991sn ; Xu:1992vc ; Cheng:1993gf ; Xu:1992sw ; Zenczykowski:1993jm ; Sharma:1998rd , current algebra Korner:1978ec ; Uppal:1994pt , and SU(3) flavor symmetry approaches Lu:2016ogy ; Savage:1989qr ; Kohara:1991ug ; Verma:1995dk ; Chau:1995gk .
-Its decay falls into three categories: Cabibbo-favored (CF) decays, singly Cabibbo-suppressed decays, and doubly Cabibbo-suppressed (DCS) decays.
+Its decays fall into three categories: Cabibbo-favored (CF) decays, singly Cabibbo-suppressed decays, and doubly Cabibbo-suppressed (DCS) decays.
 The decay amplitudes of the CF and DCS modes are expected to be proportional to the products of the Cabibbo-Kobayashi-Maskawa elements $|V_{ud}^{*}V_{cs}|$ and $|V_{us}^{*}V_{cd}|$, respectively.
@@
 In addition to direct measurements of DCS decays, the amplitudes of DCS modes can be probed using the $K^{0}_{S}$-$K^{0}_{L}$ asymmetry in the decays into neutral kaons, which arises from the interference between CF and DCS amplitudes Bigi:1994aw ; Wang:2017gxe .
-The $K^{0}_{S}$-$K^{0}_{L}$ asymmetry has been studied in the decays of charmed $D$ mesons, where the asymmetry is defined by
+The $K^{0}_{S}$-$K^{0}_{L}$ asymmetry has been studied in the decays of charmed $D$ mesons, where the asymmetry is defined as
 $R(D,K_{S,L}^{0}X)=\frac{\mathcal{B}(D\to K^{0}_{S}X)-\mathcal{B}(D\to K^{0}_{L}X)}{\mathcal{B}(D\to K^{0}_{S}X)+\mathcal{B}(D\to K^{0}_{L}X)},$ (1)
 and $X$ can be $\pi^{0}$, $\eta$, $\eta^{\prime}$, $\omega$, $\rho^{0}$ or $\phi$.
-A large asymmetry for $R(D^{0},K_{S,L}^{0}\pi^{0})$ was reported in a previous measurement by the CLEO experiment as $R(D^{0},K_{S,L}^{0}\pi^{0})=0.108\pm 0.025\pm 0.024$ CLEO:2007rhw , where the first uncertainty is statistical and the second is systematic.
+A large asymmetry of $R(D^{0},K_{S,L}^{0}\pi^{0})=0.108\pm 0.025\pm 0.024$ CLEO:2007rhw was reported in a previous measurement by the CLEO experiment, where the first uncertainty is statistical and the second is systematic.
 The BESIII experiment reported measurements of the $K^{0}_{S}$-$K^{0}_{L}$ asymmetries $R(D^{0},K_{S,L}^{0}X)$, where $X=\phi$, $\eta$, $\omega$ BESIII:2022xhe .
@@
 For the production of $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ events, the Born cross-section line shape from BESIII measurements is used BESIII:2017kqg ; BESIII:2023rwv .
-Exclusive $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ signal MC samples are generated with $\bar{\Lambda}_{c}^{-}$ decaying to twelve specific tag modes (as described in Section 3) and $\Lambda_{c}^{+}$ decaying to $pK^{0}_{L}$, $pK^{0}_{L}\pi^{+}\pi^{-}$ and $pK^{0}_{L}\pi^{0}$.
+Exclusive $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ signal MC samples are generated with $\Lambda_{c}^{+}$ decaying to $pK^{0}_{L}$, $pK^{0}_{L}\pi^{+}\pi^{-}$ and $pK^{0}_{L}\pi^{0}$ in half of the signal events, and $\bar{\Lambda}_{c}^{-}$ decaying to the signal final states for the other half.
+The remaining $\bar{\Lambda}_{c}^{-}$ and $\Lambda_{c}^{+}$ are required to decay to twelve specific tag modes, as detailed in Section 3.
 The angular distribution of the decay $\Lambda_{c}^{+}\to pK^{0}_{L}$ is modeled with decay asymmetry parameters obtained from Ref.
@@
 The $K^{0}_{L}$ is treated as a missing particle, and its four-momentum and mass are free in the kinematic fit.
-The $\chi^{2}$ of the kinematic fit for each signal mode is required to be less than the optimized value that maximizes the figure of merit $S/\sqrt{S+B}$, where $S$ and $B$ are the numbers of signal and background events from MC simulations, scaled to the data luminosity.
+The $\chi^{2}$ of the kinematic fit for each signal mode is required to be less than the optimized value that maximizes the figure of merit $S/\sqrt{S+B}$, where $S$ and $B$ are the numbers of signal and background events from the inclusive MC samples, scaled to the data luminosity.
+Here, the BF of the signal modes are assumed to be the same as the measured ones of $\Lambda_{c}^{+}\to K^{0}_{S}X$.
 The optimized requirements are $\chi^{2}<60$ for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\chi^{2}<25$ for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\chi^{2}<20$ for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$.
@@
 The model of the MC simulation is much simpler than the real detector performance, resulting in a difference between the data and MC simulation in the track parameters of the charged tracks BESIII:2012mpj .
+The correction parameters are obtained through a data-driven method using control samples of $e^{+}e^{-}\to p\bar{p}\pi^{+}\pi^{-}$ and $e^{+}e^{-}\to K^{+}K^{-}\pi^{+}\pi^{-}$.
 The helix parameters of the charged tracks are corrected, and the BFs are re-evaluated with the updated MC simulation samples.
@@
 The nominal amplitude models are then replaced by alternative ones with equivalent descriptions of the data.
-The alternative model of $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ is selected by removing the insignificant intermediate resonances.
-For $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, the amplitude fit is not stable due to the limited statistics of data.
-An alternative model is chosen with a similar fit quality to the nominal.
-The systematic uncertainties are determined to be 1.1% and 0.1% for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
+The alternative model of $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$ is selected by including other additional insignificant resonance.
+In the case of $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, the amplitude fit is not stable due to the limited statistics of data.
+The same components of resonances are employed as in the nominal case, with a different fit solution that yields a negative log-likelihood value agreeing closely with the nominal one.
+The systematic uncertainties are determined to be 1.1% and 0.9% for $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$ and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
   • Background shape.
@@
 The deviations between the Gaussian mean value and nominal BFs are assigned as systematic uncertainties.
-For the decay $\Lambda_{c}^{+}\to pK_{L}\pi^{0}$, the fit bias is found to be 0.27%, while for the other two signal modes it is negligible.
+For the decay $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, the fit bias is found to be 0.3%, while for the other two signal modes it is negligible.
 Other sources of systematic uncertainties, such as the BF of $\pi^{0}\to\gamma\gamma$, are neglected due to their negligible effects.
-Assuming that all sources of systematic uncertainties in the BFs measurements are uncorrelated, the quadratic sums of the different sources are considered as the total systematic uncertainties, which are 2.2%, 3.1%, and 2.3% for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
+Assuming that all sources of systematic uncertainties in the BFs measurements are uncorrelated, the quadratic sums of the different sources are considered as the total systematic uncertainties, which are 2.2%, 3.1%, and 2.5% for $\Lambda_{c}^{+}\to pK^{0}_{L}$, $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{+}\pi^{-}$, and $\Lambda_{c}^{+}\to pK^{0}_{L}\pi^{0}$, respectively.
 Table 4 lists all the systematic uncertainties discussed above.
@@
 Truth-match method - - 0.2
-Signal model - 1.1 0.1
+Signal model - 1.1 0.9
 Background shape 0.9 0.6 0.4
 Fit bias - - 0.3
-Total 2.2 3.1 2.3
+Total 2.2 3.1 2.5
 5 Summary